a variable plane is at a constant distant p from the origin and meets the coordinate axes at A,B,C find the locus of the centroid of the tetrahedron OABC IS

Grade:11

1 Answers

Arun
5 years ago
let lx+my+nz+d=p be the eq of the plane where p is the distance from the origin and l,m,n are the dc's
 
since the plane meets the co ordinate axes 
 
This meet the x-axis in A. Thus, y = z = 0, substituting to (1), 
lx = p, 
x = p/l. Hence,  
coordinate A = (p/l, 0, 0). 
Similarly, coordinate B and C are : 
B = (0, p/m, 0), 
C = (0, 0, p/n)
O=(0,0,0)
 now the centroid of the tedtrahedron is given by the formula:-
{(x1+x2+x3+x4)/4,(y1+y2+y3+y4)/4,(z1+z2+z3+z4)/4}
G(x, y, z) = {(p/4l + 0 + 0+0)/4 , (0 + p/4m + 0+0)/4, (0 + 0 + p/4n+0)/4}, 
G(x, y, z) = (p/4l, p/4m, p/4n)
 
Since a, b, and c are the d.c’s 
then a² + b² + c² = 1. 
But a = p/4x, b = p/4y, and c = p/4z. So : 
p²/16x² + p²/16y² + p²/16z² = 1
 
1/x² +1/y² +1/z² =16/p²